Determine the molar concentration of the anion (negative ion), if 1.2g of magnesium chloride, MgCl2(s) is dissolved in enough water to make 2.0L of solution.
1
Expert's answer
2020-09-23T06:16:39-0400

magnesium Chloride dissolves in water to form Mg2+ ions and Cl- ions in a reversible process.

MgCl2(s) "\iff" Mg2+ (aq) + 2Cl- (aq) in presence of water


mass of MgCl2 = 1.2g and the volume of solution =2.0 L


"molarity = (Moles of solute)\/ (litre volume of solution)"


calculating moles of magnesium Chloride

"mol = mass\/molecular weight"

molecular weight of MgCl2 = 24.0 +35.5 +35.5 = 95g

"moles = 1.2 \/ 95 = 0.01263"


mole ratio of MgCl2 : Cl- = 1:2


moles of chloride ions = 0.01263 * 2 =0.02526 moles of Chloride ions.


Molarity /concentration of Cl- is

"M = 0.02562 mol \/ 2.0L = 0.0126 M" Or 0.0126 mol/L